### v1

@@ 3.3 Effect of rotation
A new dedicated set of simulations was carried out to investigate the role of rotation on the collision outcome. Introducing rotations opens up a large parameter space in terms of the angular velocity and orientation of the angular momentum vectors for both the projectile and target clusters. To limit the number of simulations, we investigate only the case of head-on ($\theta$=0) collisions between same-size clusters (5000 particles each). Four different values of the collision energy were investigated ($\eta=3$, 6, 9 and 12). The angular velocity of the clusters was always set up as half of the breakup speed, and the angular momentum vector was either parallel or anti-parallel to one of the three Cartesian axes. The collision velocities are always along the x-axis. The rotational energy of the rubble piles was not included in the calculation of $\eta$, whose value is based on the center of mass kinetic energy of non-rotating analogues.
@@
Figure 8 shows the outcome of the simulations in terms of the mass of the largest fragment. Simulations with similar angular velocity configuration were consolidated in a single data point. For example, a simulation in which the projectile’s angular momentum is along the x-axis direction and the target’s momentum is along the z-axis direction would fall in the ”perpendicular spin” case, as well as projectile’s angular momentum along the y-axis and target along the x-axis and so on. In the cases in which more than one simulation is consolidated in a single data point, the error bar shows the dispersion of the various realizations. A particularly relevant case is the one of parallel spins, since planetesimals embedded in an accretion disk are expected to be spun up by the disk differential rotation with parallel angular velocities.
Figure 8: Effect of pre-collision spinning on collision outcome for head-on collisions ($\theta=0$).
@@
Our limited investigation in the effect of spin on the collision outcome reveals that anti-parallel spin have a very small effect, while both parallel and perpendicular spin make the collision more disruptive (the largest remmnants’ masses being smaller). This could be due to the fact that the surfaces of rubble piles colliding with either perpendicular or parallel spins have a strong sheer velocity, while rubble piles with anti-parallel spins would roll on each other. The latter is, however, true only because the two rubble piles have the same size and same angular velocity, in our case. A more thorough investigation of the effect of spin is needed to generalize these results.
@@ 4 Summary, discussion, and conclusions
We have presented a novel implementation of a discrete-element, soft sphere code for computing the dynamics of self-gravitating rubble piles. The code allows for particles of different sizes, includes inelastic normal forces between particles, gravity, kinetic friction, and rolling friction. Inelastic collisions are implemented in such a way to ensure constant coefficient of restitution, irrespective of particle size. Contact forces are not included at this stage, but are expected to be irrelevant at the particle dimensions investigated here. As the code is expanded to allow for simulation of small particles (e.g., dust grains) Van der Waals forces will be included (Kolanz et al., in preparation).
@@
Validation runs for the code have been successfully carried out to test its reliability by reproducing previous results from analogous, well-established codes (Figure 3; Leinhardt & Stewart 2012). Given the proven validity of the code we have carried out two experiments. In the first one we have studied the particle size composition of the largest remnant from rubble pile collisions, looking for patterns of change. We found that low-velocity collisions tend to eject predominantly small particles, resulting in remnants with larger average particle size. We note that our input rubble piles do not have a segregated particle distribution with smaller particles on the surface. For this reason, more realistic simulations may find an even larger effect. For large velocity impacts that leave behind small fragments, instead, we find that the average particle size is smaller. We ascribe this behavior to the fact that the larger particles appear to be subject to the largest amount of stress, creating the force chains that propagate the force of the impact through the rubble piles (see Figures 6 and 7). Studying the strain distribution among particles of different sizes also led us to note that the post-collision configuration of the clusters is not congruent with their initial condition. The smallest particles carry an increased fraction of the strain in the post-collision configuration. This finding emphasizes the fact that a better approach to building initial cluster configurations is by hyerarchically colliding smaller clusters rather than by a monolithic collapse as described in Sect. 2.6. We plan to investigate the effect of repeated collision in a future publication.
@@
In a second experiment we simulated collisions between fast-spinning rubble piles. We investigated the case of two identical rubble piles colliding head on ($\theta$=0). We found that collisions between rotating rubble piles lead to smaller remnants for most rotation directions. An exception was the case in which the rotation produces the effect of the two rubble piles rolling on each other. This reduced remnant size effect would be relevant in particular for collisions within accretion disks, in which the angular velocity vectors of the projectile and target objects are expected to be aligned.
@@
While DECCO is not the only code able to carry out the simulations described here, it is important to assess the robustness of numerical results by using different codes with their own intrinsic strength and weaknesses. Future plans in term of both science and development involve adding contact forces to allow for simulations of small particles, parallelizing the code to allow for a larger number of particles, and making the code GPU portable to increase efficiency.
@@ Acknowledgements.
This project was supported by NASA APRA award 80NSSC19K0330.

### v2

@@ 4.2 Effect of rotation
A new dedicated set of simulations was carried out to investigate the role of rotation on the collision outcome. Introducing rotations opens up a large parameter space in terms of the angular velocity and orientation of the angular momentum vectors for both the projectile and target clusters. To limit the number of simulations, we investigate only the case of head-on ($\theta=0^{\circ}$) collisions between same-size clusters (5000 particles each). Four different values of the collision energy were investigated ($\eta=3$, 6, 9 and 12). The angular velocity of the clusters was always set up as half of the breakup speed, and the angular momentum vector was either parallel or anti-parallel to one of the three Cartesian axes. The collision velocities are always along the x-axis. The rotational energy of the rubble piles was not included in the calculation of $\eta$, which only considers the translational kinetic energy of the cluster. At half breakout speed, the kinetic energy associated with the cluster rotation would be approximately 10% of its potential binding energy, and therefore the $\eta$ values of the results in Figure 9 would move to the right by 0.1 for all rotating cases (orange, green, and red).
Figure 8: Stress analysis among the different particle sizes for a $\theta=22^{\circ}$, $m_{p}/m_{t}=0.25$ set of collisions of increasing energy (from left to right, $\eta=0.5$, $3.0$, and $12$). The collision takes place, in all cases, at $t=0$. As detailed in the text, values larger than unity denote particle sizes that are subject to an amount of stress that is larger than the average. The figure shows that large particles always carry the largest share of the stress to counter self-gravity in the clusters, and that this imbalance is even more pronounced during high-velocity impacts. This figure includes all 6250 particles in the simulation. A similar analysis restricted to only consider particles that remain bound shows comparable qualitative behavior, but the high stress spike in the $\eta=12$ collision at $t\sim 5$ s is significantly reduced, indicating that the particles subject to the highest stress during the collision are eventually ejected.
@@
Figure 9 shows the outcome of the simulations in terms of the mass of the largest fragment. Simulations with similar angular velocity configuration were consolidated in a single data point. For example, a simulation in which the projectile’s angular momentum is along the x-axis direction and the target’s momentum is along the z-axis direction would fall in the ”perpendicular spin” case, as well as projectile’s angular momentum along the y-axis and target along the x-axis and so on. In the cases in which more than one simulation is consolidated in a single data point, the error bar shows the dispersion of the various realizations. A particularly relevant case is the one of parallel spins, since planetesimals embedded in an accretion disk are expected to be spun up by the differential rotation of the disk with parallel angular velocities.
Figure 9: Effect of pre-collision spinning on the mass of the largest fragment for head-on collisions ($\theta=0^{\circ}$). The figure shows the ratio of the mass of the largest fragment over the target mass vs. the energy parameter $\eta$ (defined in Eq. 18). Different spin configurations have been explored and are compared with non-spinning simulations (see legend).
@@
Our limited investigation in the effect of spin on the collision outcome suggests that anti-parallel spin has a very small effect, while both parallel and perpendicular spin make the collision more disruptive (the largest remnants’ masses being smaller). This could be due to the fact that the surfaces of rubble piles colliding with either perpendicular or parallel spins have a strong sheer velocity, while rubble piles with anti-parallel spins would roll on each other. The latter is, however, true only because the two rubble piles have the same size and angular velocity, in our case. A more thorough investigation of the effect of spin is needed to generalize these results. An analysis of the effect of rotation was carried out by Ballouz and co-authors (Ballouz et al., 2014, 2015), who investigated the effect of the target rotation in asymmetric collisions (the mass of the projectile being one tenth the mass of the target). Their findings are similar to ours in the fact that faster rotation is associated with smaller fragments (see, e.g., Figure 2 in Ballouz et al. 2014).
@@ 5 Summary, discussion, and conclusions
We have presented a new implementation of a soft-sphere, discrete element code for computing the dynamics of self-gravitating rubble piles. The code allows for particles of different sizes, includes inelastic normal forces between particles, gravity, kinetic friction, and rolling friction. Inelastic collisions are implemented by changing the spring constant instantaneously during rebound, ensuring constant coefficient of restitution irrespective of particle size. Short-range forces like Van der Waals forces are not included at this stage, but are expected to be irrelevant at systems size and collision velocities investigated here. As the code is expanded to allow for simulation of small particles (e.g., dust grains) these forces will be included (Kolanz et al., in preparation). Additional limitations of the code are the absence of static and twisting friction, which — together with the instantaneous change of the spring constant— limit its applicability only to dynamic calculations, and the fact that particle deformation is not accounted for in calculating surface velocities and torques. This latter simplification is not expected to significantly affect the results as the maximum compression is kept small to ensure stability (see Sect. 2.5). Post-processing analysis of the presented results has revealed that even in the fastest collisions that we have analyzed ($\eta=12$) the average spring compression is $\sim 0.03$ per cent, with less than 1 per cent of the particles ever experiencing a compression value larger than 1 per cent of their radius.
@@
Tests of energy and momentum conservation, as well as validation runs for the code have been successfully carried out to test its reliability by reproducing previous results from an analogous, well-established code (Figure 3; Leinhardt & Stewart 2012) [^8] See, however, Sect. 3.1 for a discussion on the technical and material differences between our simulations and the validation runs they are compared to.. Given the proven validity of the code we have carried out two experiments. In the first one we have studied the particle size composition of the largest remnant from rubble pile collisions, looking for patterns of change. We found that low-velocity collisions tend to predominantly eject small particles, resulting in remnants with larger average particle size. We note that our input rubble piles do not have a segregated particle distribution with smaller particles on the surface. For this reason, more realistic simulations may find an even larger effect. For large velocity impacts that leave behind small fragments, instead, we find that the average particle size is smaller. We ascribe this behavior to the fact that the larger particles appear to be subject to the largest amount of stress, creating the force chains that propagate the force of the impact through the rubble piles (see Figures 7 and 8). It should also be noted, however, that we have only tested one particle size distribution that has few large particles in it, and the role of the distribution on the collision outcome should be investigated to establish the robustness of this result. Studying the stress distribution among particles of different sizes also led us to note that the post-collision configuration of the clusters is not congruent with their initial condition. The smallest particles carry an increased fraction of the stress in the post-collision configuration. This finding emphasizes the fact that a better approach to building initial cluster configurations is by hierarchically colliding smaller clusters rather than by a monolithic collapse as described in Sect. 2.6. We plan to investigate the effect of repeated collision in a future publication.
@@
In a second experiment we simulated collisions between fast-spinning rubble piles. We investigated the case of two identical rubble piles colliding head on ($\theta=0^{\circ}$). We found that collisions between rotating rubble piles lead to smaller remnants for most rotation directions. An exception was the case in which the rotation produces the effect of the two rubble piles rolling on each other. This reduced remnant size effect would be relevant in particular for collisions within accretion disks, in which the angular velocity vectors of the projectile and target objects are expected to be aligned.
@@
While DECCO is not the only code able to carry out the simulations described here, it is important to assess the robustness of numerical results by using different codes with their own intrinsic strength and weaknesses. Future plans in term of both science and development involve adding short-range and cohesion forces to allow for simulations of small particles and microscopic systems, parallelizing the code to allow for a larger number of particles, and making the code GPU portable to increase efficiency.
@@ Acknowledgements.
We are grateful to the anonymous referees for their numerous constructive comments and suggestions that significantly improved this manuscript. This project was supported by NASA APRA award 80NSSC19K0330.